OpenStudy (anonymous):

[2.08]The formula for the perimeter P of a rectangle with length l and width w is P = 2l + 2w. Which of the following is a formula for the length of a rectangle in terms of the perimeter and width?

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

Is it multiple choice? it says which of the following... So you need to find l instead of P. Get l alone by subtracting 2w from each side. P-2w=2l. Then you can divide the 2, from the 2l, to get your answer. 1/2P-2w=l Make sense?

OpenStudy (anonymous):

yes. choces are l-(p-w)/2 l=(p+w)/2 l=(p-2w)/2 l=(p+2w)/2

OpenStudy (anonymous):

@wrstlr3232

OpenStudy (anonymous):

l= on the top not l-

OpenStudy (anonymous):

Ok, should be C. l=(p-2w)/2

OpenStudy (anonymous):

thanks!!!!

OpenStudy (anonymous):

Yep, no problem.